Morning Dove Company manufactures one model of birdbath, which is very popular. Morning Dove sells all units it produces each month. The relevant range is 0–1,900 units, and monthly production costs for the production of 1,600 units follow. Morning Dove’s utilities and maintenance costs are mixed with the fixed components shown in parentheses.  

Production Costs Total Cost
Direct materials $ 2,100
Direct labor 6,700
Utilities ($150 fixed) 570
Supervisor’s salary 3,500
Maintenance ($280 fixed) 510
Depreciation 900

Suppose it sells each birdbath for $22.

Required:

1. Calculate the unit contribution margin and contribution margin ratio for each birdbath sold.

2. Complete the contribution margin income statement assuming that Morning Dove produces and sells 1,800 units.
Answer #1

Question 1

Unit Contribution Margin = Sales Price per Unit - Variable Cost per Unit

Sales Price per Unit = 22

Variable Costs per Unit = $ 5.91

Unit Contribution Margin = 22 - 5.91

Unit Contribution Margin = $ 16.09

Notes

Calculation of Variable Cost per Unit

Direct Materials per Unit = 2100 / 1600 Units = $ 1.32 per Unit

Direct Labour Per Unit = 6,700 / 1,600 Units = $ 4.19 per Unit

Utilities Cost per Unit = 420 / 1,600 Units = $ 0.26 per Unit

Maintenance Costs per Unit = 230 / 1,600 Units = $ 0.14 per Unit

Total Variable Cost per Unit = 1.32 + 4.19 + 0.14 + 0.26 = $ 5.91

Contribution Margin Ratio = Unit Contribution Margin / Unit Sales Price * 100

Contribution Margin Ratio = 16.09 / 22 * 100

Contribution Margin Ratio = 73.14

Question 2

Contribution Margin Format Income Statement

For 1,800 Units

Particulars Amount Amount
Sales Revenue 39,600
Variable Costs
Direct Materials 2,362
Direct Labour 7,538
Utilities Cost 473
Maintenance Cost 259
Total Variable Costs 10,632
Contribution Margin 28,968
Fixed Costs
Utilities Costs 150
Supervisor Salary 3,500
Maintenance Costs 280
Depreciation 900
Total Fixed Costs 4,830
Operating Income 24,138

Notes

Contribution Margin = Sales Revenue - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Sales = 1,800 Units * $ 22 per Unit = $ 39,600

Direct Materials per Unut = 1,800 Units * $ 1.32 per Unit = $ 2,362

Direct Labour Per Unit = 1,800 Units * $ 4.19 per Unit = $ 7,538

Utilities Cost per Unit = 1,800 Units * $ .26 Per Unit = $ 473

Maintenance Cost per Unit = 1,800 Units * $ 0.14 per Unit = $ 259

Fixed Z remain same within a relevant range of production.

Additional Note

There are Normal Rounding off Adjustments being done in the value of figures so that answers are absolute figures.